OpenStudy (irastris):

What is the equation of a line that is parallel to −3x+4y=4 and passes through the point (4,0)?

OpenStudy (irastris):

Also, I need it to be in slope-intercept form.

OpenStudy (ajspeller):

The equation is in standard form so use the expression, -(A/B) to find the slope. Standard form: Ax+By=C Remember that parallel lines have the same slope. Then use y=mx+b and substitute in the slope and the point (0,4) to find b. Then use y=mx + b and only substitute in the slope and b

OpenStudy (mathmale):

There's a fast way to solve this. Take the given -3x+4y=4 and replace the final 4 with C (a constant). Now substitute the x- and y-coordinates from the given pointinto your new equation, -3+4y=C. Calculate C. Rewrite your equation with this result substitutted for C. Solve the resulting equation for y in slope-intercept form y=mx+b.

OpenStudy (mathmale):

This method works becasue 2 parallel lines have the same slope.

OpenStudy (irastris):

Thank you for the help!